OpenStudy (anonymous):

PLEASE HELP Use basic identities to simplify the expression (csc theta) (cot theta)/sec theta cot2θ csc2θ 1

OpenStudy (anonymous):

make your qu. clear.

OpenStudy (anonymous):

it is clear...

OpenStudy (zehanz):

As I am from Europe, I'm not used to sec and csc functions, so I'll just convert them back to sin and cos functions:\[\frac{ \csc \theta \cot \theta }{ \sec \theta }=\frac{ \frac{ 1 }{ \sin \theta }\frac{ \cos \theta }{ \sin \theta } }{ \frac{ 1 }{ \cos \theta } }=\frac{ \cos^2 \theta }{ \sin^2 \theta }=\cot^2 \theta\]This last answer seems short enough to me! ;)

OpenStudy (anonymous):

thank you! can you answer another quick question for me?

OpenStudy (zehanz):

Shoot!

OpenStudy (anonymous):

nevermind! i figured it out! thanks though!